OpenStudy (legends):

Wht is 80/148 in simpest form.

OpenStudy (ilovewolf):

80/148= 20/37 80/4=20 148/4=37 so then you get your 23/37 Did that make sense?

OpenStudy (legends):

@ilovewolf Why 4

OpenStudy (legends):

Is it because they both can be divided by 4

OpenStudy (triciaal):

original correct every time way to do this find the prime factors of each number and cancel the common factors multiply what is left with practice you will recognize some factors and therefore skip some steps

OpenStudy (triciaal):